Regression Analysis Estimate the beta of Amazon. 1. Use Yahoo Finance, download Amazons historical monthly stock prices for the time period (1/1/2009- 12/31/2018) and calculate monthly holding period returns. Holding period return (Ending price-Beginning Price)/ Beginning price. 2. Use Yahoo Finance, download S&P 500 historical monthly prices for the time period (1/1/2009 -12/31/2018) and calculate monthly holding period returns. Holding period return (Ending price - Beginning Price) / Beginning price. 3. Use 1% as Risk-free rate during these periods. 4. Run a regression of Y on X. The Y variable is Amazon excess return (which is Amazons return - risk free rate). X is S&P 500 excess return (or S&P 500 return - risk free rate). Show the regression summary result, the coefficient on X is an estimate of Amazons beta. Please use the rubric below to guide the development of your assignment Submit the assignment in a new Excel file. M1 Assignment-Sensitivity Analysis Rubric
Answer #1

Calculation of HPR for Amazon and S&P 500 and beta for Amazon using regression:

Risk Free rate Date AMZN S&P 500 AMZN S&P 500 AMZNS&P 500 Holding period Holding Excess Excess Closing Price Closing Price return (HPR riod return Return Return 1/1/2009 2/1/2009 3/1/2009 4/1/2009 8/1/2018 9/1/2018 10/1/2018 11/1/2018 12/1/2018 58.82 64.79 73.44 80.52 2,012.71 2,003.00 1,598.01 1,690.17 1,501.97 825.88 735.09 797.87 872.81 2,901.52 2,913.98 2,711.74 2,760.17 2,506.85 10.15% 13.35% 9.64% 13.24% -0.48% 20.22% 5.77% -11.13% -10.99% 8.54% 9.39% 3.03% 0.43% 6.94% 1.79% 9.18% 9.15% 12.35% 8.64% 12.24% -1.48% -21.22% 4.77% -12.13% -11.99% 7.54% 8.39% 2.03% -0.57% -7.94% 0.79% -10.18% 121 122 125 126 SUMMARY OUTPUT 128 Regression Statistics 131 132 133 Multiple R R Square Adjusted R Square Standard Error Observations 0.449131019 0.201718672 0.194895755 0.07658215 119 136 137 ANOVA gnificance F Regression Residual Total 1 0.173392818 0.173392818 29.56487125 117 0.686184611 0.005864826 118 0.859577429 3E-07 139 142 Coefficients Standard Erro t P-value Lower 95%U er 95 %ower 95.09 pper 95.0% Intercept X Variable 1 0.020974406 0.007020303 2.987678164 0.003425653 0.007071 0.034878 0.007071 0.034878 0.9962124980.18321625 5.437358849 3.0009E-07 0.633362 1.359063 0.633362 1.359063 145 146 147 148 Thus the coefficient is Hence Beta for Amazon isFormula sheet

A B C D E F G H I J K L
2 Risk Free rate 0.01
3 Date AMZN S&P 500 AMZN S&P 500 AMZN S&P 500
4 Closing Price Closing Price Holding period return (HPR) Holding period return (HPR) Excess Return Excess Return
5 39814 58.82 825.880005
6 39845 64.790001 735.090027 =(D6-D5)/D5 =(E6-E5)/E5 =F6-$D$2 =G6-$D$2
7 39873 73.440002 797.869995 =(D7-D6)/D6 =(E7-E6)/E6 =F7-$D$2 =G7-$D$2
8 39904 80.519997 872.809998 =(D8-D7)/D7 =(E8-E7)/E7 =F8-$D$2 =G8-$D$2
120 43313 2012.709961 2901.52002 =(D120-D119)/D119 =(E120-E119)/E119 =F120-$D$2 =G120-$D$2
121 43344 2003 2913.97998 =(D121-D120)/D120 =(E121-E120)/E120 =F121-$D$2 =G121-$D$2
122 43374 1598.01001 2711.73999 =(D122-D121)/D121 =(E122-E121)/E121 =F122-$D$2 =G122-$D$2
123 43405 1690.170044 2760.169922 =(D123-D122)/D122 =(E123-E122)/E122 =F123-$D$2 =G123-$D$2
124 43435 1501.969971 2506.850098 =(D124-D123)/D123 =(E124-E123)/E123 =F124-$D$2 =G124-$D$2
125
126
127 SUMMARY OUTPUT
128
129 Regression Statistics
130 Multiple R 0.449131019219171
131 R Square 0.201718672424851
132 Adjusted R Square 0.194895755095149
133 Standard Error 0.0765821502576298
134 Observations 119
135
136 ANOVA
137 df SS MS F Significance F
138 Regression 1 0.173392817869697 0.173392817869697 29.5648712533437 3.00089975764199E-07
139 Residual 117 0.686184611355616 0.00586482573808219
140 Total 118 0.859577429225313
141
142 Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
143 Intercept 0.0209744059937192 0.00702030300509093 2.98767816410619 0.00342565328722422 0.007071063765257 0.0348777482221814 0.007071063765257 0.0348777482221814
144 X Variable 1 0.996212498399354 0.183216250031647 5.43735884905013 3.00089975764204E-07 0.633362315984401 1.35906268081431 0.633362315984401 1.35906268081431
145
146 Thus the coefficient is =D144
147 Hence Beta for Amazon is =D146
148